Circular No. 7821

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2002 B2 (LINEAR) An apparently asteroidal object reported by LINEAR (discovery ob-

servation below) and posted on the NEO Confirmation Page has been found to be cometary by M. Tichý and J. Tichá at Kleť (coma diameter 6′′, with m_{1} = 17.2, and faint 9′′ tail in p.a. 90^{o} on Feb. 3.05 UT) and by R. H. McNaught at Siding Spring (8′′ coma and 10′′ tail in p.a. 60^{o} on Feb. 6.66).

2002 UT  α_{2000}  δ_{2000}  m_{2}
Jan. 23.39930  11^{h}40^{m}59.^{s}52  −28^{o}25′48′′.6  18.7

The available astrometry (Jan. 23–Feb. 6), the orbital elements below, and an ephemeris appear on MPEC 2002-C27.

T = 2002 Apr. 6.689 TT  ω = 257.^{o}042
Ω = 54.313 }2000.0
q = 3.84315 AU  i = 152.870

SUPERNOVA 2002at IN NGC 3720 B. Swift and W. D. Li, University of California at Berkeley, report

the LOTOSS (cf. IAUC 7514) discovery of an apparent supernova (mag ∼ 17.8) on unfiltered CCD images taken with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) on Feb. 5.4 and 6.4 UT. SN 2002at is located at α = 11^{h}32^{m}21^{s}.95, δ = +0^{o}48′08′′.8 (equinox 2000.0), which is 4′′.8 east and 6′′.0 south of the nucleus of NGC 3720. A KAIT image taken on 2001 May 6.2 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2002ap IN M74 P. Rodriguez Pascual, R. Gonzalez Riestra, B. Gonzalez Garcia, M.

Santos Lleo, M. Guainazzi, and N. Schartel, XMM-Newton Science Operations Centre, European Space Agency, write that XMM-Newton observed the location of SN 2002ap from Feb. 3.00 to 3.40 UT. SN 2002ap appears in the OM UVW1 images (wavelength range 245–320 nm) with a brightness of 14.4 counts/s, which corresponds to a flux of 7 × 10^{−15} erg cm^{−2} s^{−1} A^{−1}. Preliminary analysis of the EPIC-pn image shows a 3.5σ excess of photons that is located at a distance of 4′′ with respect to the position given on IAUC 7810. This excess corresponds to a countrate of 9.0 (± 2.5) × 10^{−4} count/s. At this stage of data reduction, the ultraviolet position error is expected to be < 10′′.

2002 February 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green